# Scaling-laws for Large Time-series Models

### v1

@@ Abstract
Scaling laws for large language models (LLMs) have provided useful guidance on how to train ever larger models for predictable performance gains. Time series forecasting shares a similar sequential structure to language, and is amenable to large-scale transformer architectures. Here we show that foundational decoder-only time series transformer models exhibit analogous scaling-behavior to LLMs, while architectural details (aspect ratio and number of heads) have a minimal effect over broad ranges. We assemble a large corpus of heterogenous time series data on which to train, and establish, for the first time, power-law scaling relations with respect to parameter count, dataset size, and training compute, spanning five orders of magnitude.
@@ 1 Introduction
The emergence of large language (LLMs; [9, 10, 11, 12]) and computer vision models [13, 14, 15, 16, 17, 18, 19] with zero-shot prediction capabilities has sparked a growing interest in developing foundation models for time-series — general purpose models for forecasting, pre-trained on a large and diverse corpus of time-series data, with the aim of achieving state-of-the-art zero-shot forecasting performance across many domain areas [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Large time-series models (LTMs) are already achieving zero-shot prediction capability that is similar to (and in some cases better than) baseline statistical or domain-specific models in many areas [20, 21, 22, 23, 24, 25, 26, 27, 28, 34].
@@
Underpinning the resource investment and subsequent success of LLMs and large-scale computer vision models was the demonstration of neural scaling laws [35, 36, 37, 38, 39, 18, 40]. The observed power-law scaling of test loss with model sizes, compute resources, and training set sizes, has provided a basis for predicting the expected gains from different efforts, aiding the community in allocating resources appropriately to achieve breakthroughs in performance. The establishment of similarly favourable neural scaling laws for LTMs would serve as both a motivation and guide in the pursuit of foundation models for time-series forecasting.
@@
Contributions: In this work, we establish for the first time neural scaling laws for large time-series models (LTMs), demonstrating that LTMs enjoy similar power-law scaling laws to language and computer vision. We train decoder-only transformer models (with architectures tailored to time-series forecasting; §3.1) on a large, diverse, and well-balanced dataset comprising around 8 billion data points across 30,211,687 individual time-series, drawn from $38$ qualitatively distinct data sources from varied areas (see §2). We demonstrate power-law like scaling behavior of model performance with model size, compute, and dataset size (Fig. 1). For each scaling case-study, we show similar scaling behavior in three key measures of the model performance: the mean-square error (MSE) characterizing the accuracy of point (posterior mean) forecasts; the Continuous Ranked Probability Score (CRPS [41]) characterizing the fidelity of the probabilistic predictions (ie., coverage of the forecast posterior density); and the log-likelihood loss characterizing the Kullback-Leibler (KL) divergence between the model and data generative distributions.
Summary of Results: In Fig. 1 we show the scaling of model performance (MSE, CRPS and log-likelihood) with model size (top row), compute (middle row), and dataset size (bottom row). We observe power-law scaling in the log-likelihood across the full range of parameters tested, whereas the MSE and CRPS show broken power-law behavior. Parameters of the fitted power-law functions can be found in Tab. 2. In Fig. 3 we show that the model performance is only weakly sensitive to architecture details (aspect ratio, and the number of heads).
Compute Requirements: To produce the results in this paper requires $\mathcal{O}(50-70)$ individual production runs. Apart from the 100M parameter run, these were all carried out on single A100 NVIDIA GPU instances, each taking between 1 and 3 days to complete. As such, overall, the work presented here required $\mathcal{O}(150)$ GPU-days of compute. To host the full dataset, we also required a CPU RAM allocation of approximately 250 GB.
Refer to captionRefer to captionRefer to caption
Figure 1: Test Loss Scaling Laws: Here we show three metrics for the minimum in-sequence test loss as a function of the number of parameters, compute, and dataset size. We demonstrate that, like language, the performance of large time-series models scales approximately as a power law with all three quantities. Note that we use the log-likelihood as the actual loss function for training. To ensure it remains positive for the power-law scaling we add a small constant.
2 Data
The development of a foundation model for time-series forecasting is predicated on the availability of a sufficiently large, diverse, and well-balanced dataset to train on. We constructed a corpus of time-series data comprising around $8$ billion data points drawn from $38$ varied data sources (see Tab. 1). For the purpose of this study, our focus was to ensure our dataset is: large enough so that for our largest models ($\sim 100$M parameters) we are still operating in the approximately infinite data limit (as discussed in the original LLM scaling-law paper [35]); as diverse as possible given the practical limitations on publicly available data; and balanced, such that no individual dataset comprises more than roughly $15\%$ of the total number of data points. Our resulting dataset is competitive with the state-of-the-art in terms of both diversity and size [^1] Note that where other state-of-the-art time-series datasets from recent studies are larger, this discrepency is largely accounted for by either the use of synthetic data in those studies (which we do not include in this work), and/or the reduction in our total data point numbers due to re-balancing of our data to ensure it is not dominated by a single data source., while covering a wide variety of sampling frequencies, record lengths, dynamic ranges, and underlying latent process phenomenology. In this study we focus exclusively on univariate time-series forecasting, and leave the study of scaling-laws for multivariate LTMs to future work.
The data sources used to construct our dataset are summarized as follows (see also Tabs. 1, 37):
Monash: The Monash repository [42] is an open-source collection of data from many sources. We exclude datasets that are too short (shorter than our context length of $256$; see § 3.1), and removed individual time-series that we found to destabilise training (due to e.g., a combination of heavy-tailed behavior and a large fraction of missing data). The included Monash data is summarized in Tab. 3.
Climate: We include climate data from two sources: the National Oceanic and Atmospheric Administration (NOAA), and the fifth generation European Centre for Medium-Range Weather Forecasts atmospheric reanalysis of the global climate (ERA5). From these we select a variety of different observables and sampling frequencies, summarized in Tabs. 45.
Energy: Here we use a subset of the data from the BuildingsBench dataset [43]. The data is sampled at a fixed 15-minute cadence and represents energy demand across US commercial and real estate buildings.
Traffic: We use the LargeST [44] traffic dataset from 2017-2021 which represents traffic flow in California and is made up (once processed) of a total of 8520 series.
Finance: We source daily returns and trading volume data for $5038$ stocks lists across various exchanges from yahoo finance (see Tab. 6). For trading volume, we model the logarithm of the data (due to the large dynamic range).
Audio: We include a diverse set of audio data from the command, bird, and Arabic speech datasets [45, 46]; see Tab. 7.
@@ 2 Data
3 Methods
3.1 Model and Training Details
Decoder-only Transformer: We use decoder-only transformer models with self-attention as the primary architecture throughout, with a learned positional-encoding and flexible distribution head to make it more amenable to probabilistic time-series forecasting (as discussed below). We use a context length of $256$ data points, and ReLU activation functions throughout.
Learned Positional Encoding and Embedding: Following the performance gains shown in Ref. [26], we use a learnable encoding rather than the sinusoidal positional encoding used in the original transformer model [47]. Both the learned positional encoding and embedding are simple linear layers going from one input to $d_{\mathrm{m}}$ outputs (see below). The positional encoding is treated in the usual fashion by just adding elementwise to the embedded float.
Distribution Head: The distribution head defines the distribution family modelled by the transformer, and is a critical architecture choice in enabling probabilistic forecasts with accurate coverage, as well as stable training and convergence. We use a Student’s-$t$ distribution head, where the mean, $\mu$, variance, $\sigma^{2}$, and degree of freedom, $\nu$, are separate outputs of the model. [^3] We apply softplus activations to the distribution head outputs corresponding the variance and degrees-of-freedom, to ensure positive definiteness. The Student’s-$t$ distribution head allows us to model heavy-tailed data and processes, and in experiments we found that the Student’s-$t$ head enabled significantly more stable training than a Gaussian head (or simple posterior mean forecasting with an MSE loss). We note that in reality, many time-series data and underlying processes exhibit skewness, and a distribution head family that can describe asymmetric distributions would be well-motivated (e.g., Generalized gaussians, SinhArcSinh distributions, or more expressive normalizing flow or diffusion-based heads). We leave the exploration of more expressive distribution heads to future work.
We use a separate multi-layer perceptron (MLP) with four hidden layers of dimension $d_{\mathrm{m}}$ for each output of the distribution head, i.e., the mean, variance and degrees-of-freedom of the Student’s-$t$ distribution. In experiments we found that including separate MLPs for each distribution parameter (rather than a single MLP with three outputs), significantly stabilizes training. We use a negative log-likelihood loss throughout.
Parameter Counting: With this setup, the model architecture is defined by the following parameters: the number of output dimensions $\theta_{\mathrm{out}}$, the input/output size of the linear layers in the self-attention $d_{\mathrm{m}}$, the number of heads444 Note that adding heads does not add to the parameter count. Instead the sequence is split and passed though different heads and recombined at the end. We therefore require that $d_{\mathrm{m}}\,\mathrm{mod}\,N_{\mathrm{heads}}=0$ $N_{\mathrm{heads}}$, the hidden layer size of the linear layers directly after the self-attention $d_{\mathrm{ff}}$, and the number of decoder layers $N_{\mathrm{l}}$. Throughout this work we fix $d_{\mathrm{m}}=d_{\mathrm{ff}}$ and treat all trainable parameters (including weights and biases of all layers) equally in the parameter counting. As shown in Fig. 1, we explore models with $\sim 10^{3}$ to $\sim 10^{8}$ trainable parameters.
Training Details: We use the AdamW optimizer with a batch size of 512, a cosine learning rate scheduler with a linear warm up of 3000 training steps, and train for a total of $10^{5}$ steps. When training on the entire dataset ($\sim 8$B data points), this equates to roughly two epochs. To reduce computational costs we compute the test loss every $\mathcal{O}(200)$ steps and average over a random 10% of the test data each time.
3.2 Learning Rate and Architecture Dependence
In order to extract reliable scaling laws, we first need to determine the sensitivity of LTM performance as a function of both architecture choices (e.g., aspect ratio and number of heads), and learning rate scheduling.
Learning Rate Scheduling: In training we use a linear warm up followed by sinusoidal decay for the learning rate scheduling. In Fig. 2 we show the effect of changing the maximum learning rate reached at the end of the warm up. The performance of the model (CRPS) clearly depends on the maximum learning rate, and that dependence is itself a function of parameter count. The dependence on maxmimum learning rate is strong enough that it is possible to get better performance with a smaller model, if the maximum learning rate is too small (or too large) for the larger model. Moreover, for a fixed model size we see a clear optimum learning rate above which models diverge (shown as crosses on Fig. 2). To ensure that we used an optimal maximum learning rate as a function of model size, we fit a power law with a constant offset to the best models (at each parameter size) shown in Fig. 2. In the few cases for the largest models where our power law fit overestimates the optimal maximum learning rate (leading to divergence), we slowly reduce the learning rate until we achieve convergence.
Aspect Ratio and Number of Heads: Figure 3 shows how the minimum CRPS varies as a function of aspect ratio $d_{m}/N_{\mathrm{l}}$ (left panel) and the number of attention heads, $N_{\mathrm{heads}}$ (right panel). Overall, we find that neither of these choices have a large effect on performance; we note that this is similar to the weak architecture sensitivity observed for LLMs [35]. This is most clearly seen in the right panel of Fig. 3 where we are able to keep the number of parameters constant whilst varying the number of attention heads. For both model sizes tested, there is no significant change in the minimum CRPS achieved as more heads are added. For the main parameter, compute, and data scaling runs (the main results of this paper), we fix the number of heads to four.
Performance as a function of aspect ratio shows a slightly more complex behavior. In particular, we find smaller aspect ratios tend to perform better. However, this dependence is only weak up until aspect ratios of $>128$, after which the model performance decreases substantially. Based on these observations, we keep the aspect ratio $<70$ for all scaling runs.
4 Scaling and Optimality
We now discuss the scaling results as a function of parameter count $N_{p}$, dataset size $\mathcal{D}$, and compute $\mathcal{C}$. For each scaling-relation, we fit a power law with the form:
$L(A)=\left(\frac{A}{A_{0}}\right)^{-B_{0}}\,,$ (1)
where $L$ is the objective function (i.e., MSE, CRPS, or log-likelihood) and $A$ is the quantity being scaled (i.e., parameter count $N_{p}$, dataset size $\mathcal{D}$, or compute $\mathcal{C}$. The fitted values of the normalization $A_{0}$ and index $B_{0}$ for each scaling-law are given in Tab. 2. Where broken power-law like scaling is observed, we report the power law fit after the break only, since this is the relevant quantity to motivate extrapolation to larger models / datasets / compute resources.
4.1 Parameter Scaling
First we look at parameter scaling which is shown in the top panels of Fig. 1. In particular, we plot the in-sequence test loss computed for the MSE, CRPS, and log-likelihood (which is used as the loss function during training). For each model size we only plot the minimum test loss. We can see that all loss functions follow approximately power-law behavior over nearly five orders of magnitude in model size. This trend is particularly well observed in the log-likelihood. Note that we added a constant factor of two to the log-likelihood, to keep it positive.555 We note here that a constant additive factor can change the slope of the fitted power-law. To remain as agnostic as possible we choose to add the smallest integer required to make all values of the loss positive (i.e., two).
Interestingly, we do see a slight break in the power-law behavior in both the MSE and CRPS test losses, although this is mainly due to the behavior of models at very low parameter count. We plan to investigate the details of this break in future work. To ensure that power-law scaling isn’t biased by this break, we only fit to models with $>10^{5}$ parameters. This difference between the MSE and log-likelihood is not surprising as the MSE only tests the performance of the model to predict the mean whereas the log-likelihood captures the distribution. On the other hand, the difference in the scaling behavior seen across the CRPS score is slightly more surprising. A possible explanation for this discrepancy is that the log-likelihood is known to be more sensitive to variations in the tails of the forecast (Student’s-$t$) distribution [48]. Regardless of these details, this observed scaling over many orders of magnitude is remarkable and demonstrates that LTMs are likely to reach state-of-the-art performance given enough data and model size.
Table 2: Power-law fits.
MSE CRPS Log-Likelihood
$\log_{10}(A_{0})$ $B_{0}$ $\log_{10}(A_{0})$ $B_{0}$ $\log_{10}(A_{0})$ $B_{0}$
Number of Parameters, $N_{p}$19.47 0.042 -22.64 0.036 4.33 0.151
Training Compute, $\mathcal{C}$38.88 0.031 -43.03 0.028 -6.65 0.101
Dataset Size, $\mathcal{D}$8.91 0.062 -30.42 0.027 7.00 0.188
4.2 Data Scaling
Scaling the dataset in a coherent manner requires some care. In particular, we find that to achieve reliable scaling behavior as the dataset size grows requires one to keep the diversity approximately fixed i.e., the proportion of each dataset’s contribution to the total must remain the same (see Tab. 1). For data that is significantly longer than our context length, we simply use a random fraction, $f_{d}$, of each time series. For data that would become shorter than our context length once cut, we instead randomly drop the entire series with probability equal to $1-f_{d}$. To compute the test loss, however, we average over the full test set as in the parameter scaling runs. This is to allow direct comparison between data scaling runs with different values of $f_{d}$, and reduce the noise in the test loss computation for small dataset sizes.
Results are shown in the bottom panel of Fig. 1 where we train a $\sim 20$M parameter model using the optimum max learning rate found during the parameter scaling exploration as well as early stopping. We again find power-law scaling across approximately four orders of magnitude. Interestingly, all three loss functions scale as a power law over the full range tested, although there is a mild indication of saturation for the largest dataset sizes.666 Ideally we would use a larger model to ensure we are operating in the approximately infinite parameter limit, but we were limited by computational resources. Results of the power-law fits are shown in the bottom row of Tab. 2.
4.3 Compute Scaling
Finally, we turn to compute scaling where we again follow a similar method to Ref. [35]. In particular, the compute at any given stage in the training process is given by $\mathcal{C}=6BN_{\mathrm{p}}L_{\mathrm{seq}}$ where $B$ is the batch size, $N_{\mathrm{p}}$ is the number of parameters in the model, and $L_{\mathrm{seq}}$ is the context length. Test losses as a function of compute are shown in Fig. 1 (middle panel) and the scaling law is obtained by taking the minimum value at any given value of compute. Although we see a significant amount of noise in the loss functions during training, there is a clear overall trend towards lower test losses for higher compute which is accurately tracked by a power law.
Similarly to the parameter scaling we see a slight break at low values of compute for both the MSE and CRPS test losses. We therefore again choose to only fit to data above $10^{-7}$ PF-days. Finally, we note that the steepness of the log-likelihood test losses at large compute values indicates that the models are not yet fully converged. The effect is more dramatic at larger model sizes, indicating that further tuning of the models and learning rate scheduling can still improve performance. We empirically observe that the test loss tends to plateau in the middle of training and then decrease again as the scheduler reaches low values of the learning rate (towards the end of the sinusoidal decay). Further work is needed to test whether an alternative scheduler can improve overall performance and thus further improve model scaling.
5 Discussion and Limitations
Throughout this paper we have focused on evaluating models by their in-sequence test loss, and have not discussed explicitly how this translates into a model’s ability to forecast into the future. Again taking inspiration from LLMs, we have implicitly assumed that a model which produces good in-sequence predictions should naturally be able to forecast into the future. This is theoretically and empirically well-motivated; as the modelled posterior predictive distribution for the next value becomes increasingly well-calibrated, any accumulating errors from auto-regressively rolling out those predictions into the future should also decrease (leading to more accurate prediction further into the future). In App. B, we show some clear examples of how forecast roll-out becomes increasingly coherent with increasing model size (and corresponding decreases in test loss) [25, 34]. We leave the study of scaling-laws based on rolled out forecasting ability (on different time horizons) to future work.
We have detailed the specific scaling laws for a decoder-only transformer with self-attention. However, it would be interesting to explore how modifications to this architecture might improve model scaling. In particular, much of the recent progress in using LTMs [20, 21, 22, 23, 24, 25, 26, 27, 28, 34] has involved various changes to transformer architectures to make them more suited to time-series data. We advocate for comparative scaling law studies as new architectures are introduced, to allow the community to evaluate which model architectures will eventually reach state-of-the-art zero-shot prediction capabilities.
When experimenting with data scaling, we found it was neccessary to scale the training data in such a way as to (approximately) preserve the data diversity in the scaled training set (i.e., keep relative contributions of each data source constant). Approaches to data scaling that did not preserve data diversity failed to reveal clear scaling behaviour; we attributed this to the resulting variation in data diversity dominating over any scaling behaviour. Given the importance of training data diversity in establishing data scaling laws, and in training state-of-the-art pre-trained foundation models in general, developing a robust framework for measuring data diversity would be of great utility to the field. Taking inspiration from Ref. [49], one avenue could be looking at the impact of new data sources on gradients during training; data which produce gradients with a high cosine similarity compared to the current training set would be likely to increase data diversity. We leave exploration of measures of data diversity to future work.
An additional scaling law that we have not explored in this work (due to computational limitations) is performance as a function of increasing context length. Multiple studies (e.g., [34, 50]), both for LLMs and LTMs, have shown that increasing the context length window significantly improves both in-sequence prediction as well as forecasting. This should be particularly apparent for time series data as increasing context length gives the LTMs access to lower frequency modes of variability, which might otherwise be missed. We will explore context-length scaling in future work.
In this work we have focused exclusively on univariate time-series data. However, a general purpose foundation model for time-series forecasting should be able to cope with the more general setting of multivariate time-series prediction, while accounting for multiple exogeneous covariates. Establishing scaling-laws for multivariate time-series forecasting would be an important extension to this work. This will demand the assembly of a large and diverse training set of multivariate data, each with their own (varying number of) exogeneous factors.
Finally, the literature on transformers for time series forecasting has, to date, mostly focused on trying to achieve state-of-the-art performance on a small set of benchmark experiments. Our results indicate that large scale pre-training leads to scalable performance improvements that, given enough time and compute, can result in state-of-the-art performance across the majority of benchmarks.
Impact Statement and Ethics: The main goal of this paper is to advance the fields of machine learning and foundational time series forecasting. Time series forecasting has numerous applications across both commercial sectors and other critical areas such as healthcare, climate science, finance, and energy management. By improving forecasting techniques, our work can contribute to more accurate predictions and better decision-making in these diverse fields. Due the exploratory nature of this work, it is highly unlikely that there are direct negative societal impacts of the current study. In addition, there are no specific ethical implications of our work that are relevant for further discussion here. Indeed, we confirm that everything in this paper complies with the NeurIPS Code of Ethics.
Acknowledgments: TE thanks Jared Kaplan for providing initial motivation that these scaling laws would appear. TE is supported by the Horizon Postdoctoral Fellowship. We thank Hiranya Peiris for providing access to GPU compute resources. JA has been supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programmes (grant agreement no. 101018897 CosmicExplorer). This work partly was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the National Science Foundation (NSF) grant number OAC 1920103. Some of this work was also carried out on the Snellius Compute Cluster at SURFsara.
Appendix A Dataset Details
In this section we detail the various sources that form the basis of our dataset and the choices made during its construction. Throughout data acquisition we had two critical features that the dataset needed to satisfy to facilitate the study of the main paper. Firstly, we needed a large enough dataset to train models up to $\mathcal{O}(100)\mathrm{M}$ parameters. This is particularly important as the original work [35] demonstrated scaling laws over many orders of magnitude. Although we were limited here by computational resources, it was important to explore the parameter range $10^{3}-10^{8}$ (the current range used in other transformer based time series models). Secondly, the dataset needed to be sufficiently balanced across sources and features such that the scaling laws observed here can be considered a feature of foundation models and not associated with a particular dataset.
@@ Appendix A Dataset Details
Taking inspiration from large language models [35], we therefore aimed to gather around $\mathcal{O}(10^{10})$ data points from a variety of domains. We note, however, that treating a single floating point number on a similar footing to a language token is not necessarily a good comparison. In particular, language tokens can contain significantly more semantic meaning than a floating point number can. It’s therefore likely that we will need more data points overall, although more work is required to make this comparison more precise. Before detailing our particular sources, we would like to emphasize that there is an enormous corpus of time series data publicly available but is not currently formatted for easy downloading and processing. Ref. [25] is the first paper to open-source a fairly large dataset but significant work is still required to expand its size and diversity.777 Note that by the time this data became open-source we had already fixed our dataset. For example, large-scale, state-of-the-art language models are likely trained on well over a trillion tokens.
We now discuss each dataset presented in Tab. 1. Overall, we assembled approximately 8B data points from six domains, each with a its own source of variation.

### v2

@@ Abstract
Scaling laws for large language models (LLMs) have provided useful guidance in training ever larger models for predictable performance gains. Time series forecasting shares a similar sequential structure to language, and is amenable to large-scale transformer architectures. Here we show that foundational decoder-only time series transformer models exhibit analogous scaling-behavior to LLMs, with architectural details (aspect ratio and number of heads) having a minimal effect over broad ranges. We assemble a large corpus of heterogenous time series data on which to train, and establish for the first time power-law scaling with parameter count, dataset size, and training compute, spanning five orders of magnitude.
@@ 1 Introduction
The emergence of large language (LLMs; [9, 10, 11, 12]) and computer vision models [13, 14, 15, 16, 17, 18, 19] with zero-shot prediction capabilities has sparked an interest in developing foundation models for time-series — general purpose forecasting models, pre-trained on a large and diverse corpus of time-series data, aimed at achieving state-of-the-art (SOTA) zero-shot forecasting performance across many domain areas [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Large time-series models (LTMs) are already achieving zero-shot prediction capability similar to or better than baseline statistical or domain-specific models in many areas [20, 21, 22, 23, 24, 25, 26, 27, 28, 34].
@@
Underpinning the investment and subsequent success of LLMs and large-scale computer vision models was the demonstration of neural scaling laws [36, 37, 38, 39, 40, 18, 41]. The observed power-law scaling of test loss with model size, compute resources, and training set size, has provided a basis for predicting expected gains from different efforts, aiding the community in allocating resources appropriately to achieve performance breakthroughs. Given qualitative differences in data and modeling challenges, existence of neural scaling-laws for time-series is not guaranteed from the language and computer vision results. Establishment of similarly favourable scaling laws for LTMs would serve as a motivation and guide in the pursuit of foundation models for time-series forecasting.
@@
Contributions: We establish for the first time that LTMs enjoy similar power-law scaling laws to language and computer vision. We train decoder-only transformer models (with architectures tailored to time-series forecasting; §2) on a large, diverse, and well-balanced dataset comprising around 8 billion data points across 30,211,687 individual time-series, drawn from $38$ qualitatively distinct data sources from varied areas (see §2). We demonstrate power-law like scaling behavior of model performance with model size, compute, and dataset size (Fig. 1), finding similar scaling behavior in three key measures of model performance: the mean-square error (MSE) characterizing the accuracy of point (posterior mean) forecasts; the Continuous Ranked Probability Score (CRPS [42]) characterizing the fidelity of the probabilistic predictions (ie., coverage of the forecast posterior density); and the log-likelihood loss characterizing the Kullback-Leibler (KL) divergence between the model and data generative distributions.
2 Data and Methods
[…]
Model Architecture: We use decoder-only transformer models with self-attention as the primary architecture throughout, with a context length of $256$ data points and ReLU activation functions. Following the performance gains shown in Ref. [26], we use a learnable encoding rather than the sinusoidal positional encoding used in the original transformer model [43]. Both the learned positional encoding and embedding are simple linear layers going from one input to $d_{\mathrm{m}}$ outputs.
Distribution Head: We use a Student’s-$t$ distribution head, where the mean, variance, and degrees of freedom are each modelled by a separate dense network with four hidden layers of dimension $d_{\mathrm{m}}$. The Student’s-$t$ distribution allows us to model heavy-tailed data, which we find in experiments enables significantly more stable training than a Gaussian head or MSE loss. We note that in reality, time-series data and processes exhibit diverse distributional characteristics, and a more expressive distribution head (e.g., mixture model, normalizing flow or diffusion model) is well-motivated. We leave the exploration of scaling-laws under more expressive distribution heads to future work. We use a negative log-likelihood (KL) loss for training throughout.
Parameter Counting: With this setup, the model architecture is defined by the following parameters: the number of output dimensions $\theta_{\mathrm{out}}$, the input/output size of the linear layers in the self-attention $d_{\mathrm{m}}$, the number of heads $N_{\mathrm{heads}}$, the hidden layer size of the linear layers directly after the self-attention $d_{\mathrm{ff}}$, and the number of decoder layers $N_{\mathrm{l}}$. Throughout this work we fix $d_{\mathrm{m}}=d_{\mathrm{ff}}$ and treat all trainable parameters (including weights and biases of all layers) equally in the parameter counting. As shown in Fig. 1, we explore models with $\sim 10^{3}$ to $\sim 10^{8}$ trainable parameters.
Learning rate and architecture sensitivity: To extract reliable scaling laws, we need to determine sensitivity and robustness to the learning rate (LR) schedule and architecture choices. We use a linear warm up followed by sinusoidal decay for the learning rate scheduling, and find that the model performance clearly depends on the maximum LR reached at the end of the warm-up. To ensure robustness to the maximum LR, we fit a power-law to the best model at each parameter size to estimate the optimal maximum LR as a function of parameter count, shown in Fig. 3 (Appendix C).
[…]
3 Results
Figure 1: Test Loss Scaling Laws: Minimum MSE (left), CRPS (middle) and log-likelihood (right) in-sequence test metrics as a function of the number of parameters (top), compute (middle), and dataset size (bottom).
Scaling as a function of parameter count $N_{p}$, dataset size $\mathcal{D}$, and compute $\mathcal{C}$ is summarized in Fig.1. For each scaling-relation, we fit a power law of the form $\ln L(A)=-B_{0}\ln A+B_{0}\ln A_{0}$, where $L$ is the objective function (MSE, CRPS, or log-likelihood) and $A$ is the scaled quantity (i.e., parameter count, dataset size, or compute). The fitted parameter values are given in Tab. 7 (Appendix D). Where broken power-law like scaling is observed, we report the power law fit after the break only, since this is the relevant quantity to motivate extrapolation to larger models / datasets / compute resources.
[…]
Data Scaling: Extracting reliable scaling behavior with dataset size requires keeping the data-diversity fixed, i.e., each dataset’s relative contribution to the total data count should remain the same under scaling (see Tab. 1). For time-series that are significantly longer than our context length, we use a randomly chosen portion, $f_{d}$, of each time-series, while for series that would become shorter than our context length once cut, we instead randomly drop the entire series with probability equal to $1-f_{d}$. We compute the test loss over the full test set to allowing direct comparison between runs with different values of $f_{d}$, and reduce the noise on the test loss in the small (scaled) dataset limit.
Results are shown in the bottom row of Fig. 1 where we train a $\sim 20$M parameter model using the optimum max learning rate found during the parameter scaling exploration, with early stopping. We find power-law scaling across four orders of magnitude in all three performance measures.
Compute Scaling: The compute at any given stage in the training process is given by $\mathcal{C}=6BN_{\mathrm{p}}L_{\mathrm{seq}}$, where $B$ is the batch size, $N_{\mathrm{p}}$ is the number of parameters in the model, and $L_{\mathrm{seq}}$ is the context length [36]. Test losses as a function of compute are shown in Fig. 1 (middle row), where the scaling law is obtained from the minimum test loss attained at any given value of $\mathcal{C}$. Although we see a significant amount of noise in the loss functions during training, there is a clear overall trend towards lower test losses for higher compute, which is well-described by a power law. Similarly to the parameter scaling we see a mild break at low compute values for both the MSE and CRPS test losses. Note that while the MSE and CRPS metrics appear to be approximately converged over the compute range considered, the log-likelihood may not be fully converged; additional training may be needed to obtain accurate compute scaling-law fits for the log-likelihood in the large model limit.
4 Discussion
We have focused on evaluating models by their in-sequence (next step) test loss, rather than explicitly assessing model’s ability to forecast further into the future. The implicit assumption that a model with good in-sequence predictions should naturally be able to forecast into the future is theoretically and empirically well-motivated: as the modelled posterior predictive distribution for the next value improves, any accumulated errors from auto-regressive roll-out those predictions into the future should also improve. In App. E, we show some clear examples of how forecast roll-out becomes increasingly coherent with increasing model size. We leave the study of scaling-laws based on forecasting ability on different time horizons to future work.
We have detailed the specific scaling laws for a decoder-only transformer with self-attention. However, it would be interesting to explore how modifications to this architecture might improve model scaling. In particular, much of the recent progress in using LTMs [20, 21, 22, 23, 24, 25, 26, 27, 28, 34] has involved various changes to transformer architectures to make them more suited to time-series data. We advocate for comparative scaling law studies as new architectures are introduced, to allow the community to evaluate which model architectures will eventually reach SOTA zero-shot prediction capabilities.
When experimenting with data scaling, we found it was critical to scale the training data in such a way as to preserve the data diversity; approaches to data scaling that did not preserve data diversity failed to reveal any clear scaling behaviour. Given the importance of data diversity in establishing data scaling laws, and in training SOTA pre-trained foundation models in general, developing a robust framework for quantifying data diversity would be of great utility to the field.
One scaling law that we have not explored in this work (due to computational limitations) is performance as a function of increasing context length. Multiple studies (e.g., [34, 45]), both for LLMs and LTMs, have shown that increasing the context length significantly improves both in-sequence prediction and forecasting, and a recent study [46] find interesting scaling behaviour of LTMs with context length. We will explore context-length scaling in future work.
We have focused on univariate time-series data. However, a general purpose foundation model for time-series forecasting should be able to cope with the more general setting of multivariate time-series prediction, with multiple exogeneous covariates. Establishing scaling-laws for multivariate time-series forecasting will be an important extension to this work; this demands the assembly of a large and diverse training set of multivariate data, each with their own exogeneous factors.
Appendix A Dataset Details
@@
[…]
@@
Taking inspiration from large language models [36], we therefore aimed to gather around $\mathcal{O}(10^{10})$ data points from a variety of domains. We note that treating a single floating point number on a similar footing to a language token is not necessarily a good comparison; language tokens can contain significantly more semantic meaning than a floating point number can. The continual growth of open-source time-series datasets in both size and diversity will enable increasingly robust neural scaling studies.
Before detailing our particular sources, we would like to emphasize that there is a large corpus of time series data publicly available but is not currently formatted for easy downloading and processing. Ref. [25] was the first paper to open-source a large dataset333 Note that by the time this data became open-source we had already fixed our dataset for the production runs completed for this study. , setting a trend for improved training and benchmarking of foundational time-series models. However, significant work is still required to expand available datasets in size and diversity to reach the same maturity as LLMs (large-scale, SOTA language models are trained on well over a trillion tokens).
We now discuss each dataset presented in Tab. 1.